OpenStudy (anonymous):

Which answer shows this equation in standard form? 6 – 2(x – y) = –3x + 5 A. x + 2y = –1 B. x + 2y = 1 C. x – 2y = 1 D. –5x + 2y = –1

OpenStudy (anonymous):

A

OpenStudy (anonymous):

please break it down

OpenStudy (kamibug):

Distribute the 2 and combine like terms. :) Distribute. 6 - 2x + 2y = -3x + 5 Add 3x to both sides. 6 + x + 2y = 5 Subtract 6 from both sides. x + 2y = -1

OpenStudy (kamibug):

\(\Huge\bf\color{indigo}{Welcome~to~Openstudy!}\)

OpenStudy (anonymous):

thanks